OpenStudy (anonymous):

(1/3m+6)cube of binomial

OpenStudy (ikram002p):

u mean \(\large (\frac{ 1}{3m+6})^3\) ?

OpenStudy (anonymous):

or \(\left(\frac{1}{3}m+6\right)^3\)

OpenStudy (rational):

the first one is not a binomial, so...